Mallory sends 50,000 phishing emails., posing as an agent for Bob's Bank. 10% of the e-mails reach customers of Bob's Bank. 0.47% of those customers fall for the phishing email. Mallory collects their banking info and fradulently steals an avg of $1000 from each of them. How much money does Mallory make?
1
Expert's answer
2013-04-16T10:37:43-0400

10% of the phishing emails = 50 000 * 0.1 = 5000
0.47%оf the 5000 = 0.0047 * 5000 = 23.5
1000$from each person = 1000$ * 23.5 = 23500$
ANSWER is 23500$